Quick Start Inc. is expected to pay a dividend of $1.05 next year and dividends are expected to continue their 7% annual growth rate. The SML has been estimated as follows:

kj = 0.08 + 0.064βj

Assuming Quick Start has a beta of 1.1, what would happen to its stock price if inflation expectations went from the current 5% to 8%?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text General ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard FZ169 FS Font Alignment Number Styles Cells FT FU FV FW FX FY FZ GA GB GC GD GE GF 0.08 + 0.064 Bj Rf = 0.08 = 896, Rm-Rf= RISK PREMIUM = 6.4% 0.08+0.064( 1.1)-15.04% 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 169 170 4 KE CAPM UTILITY, SHARPE, beta bond PO = 1.05/(0.1 504-007)- 13.06 NOW IF INFLATION RATE INCREASES FROM 5 TO 8% INCREASE OF 3 % Rf WILL INCREASE BY 3%, Rf= 11% Rm-Rf= RISK PREMIUM WONT CHANGE = 6.4% NOW kj = 0.11 +0.064 Bj 0.11+0.064( 1.1)-18.04% Po - D1/(kj-g) PO = 1.05/(0.1804-007)- 9.51 ANSWER STOCK PRICE DECREASE FROM 13.06 TO 9.51 future INDEX INTL CAP BUDLEASING PV, FV, ANNUITYDIR cleanYIELD bond tru WACC RESI ex d 福 130% 18:32 15-01-2019